# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ 3.1 Dissipative channels
We aim to derive the energy-differential cross-section within a perturbative framework. For each dSIDM model, the emission amplitude $\mathcal{M}$ is constructed from Feynman diagrams as shown in Fig. 1. In the following, we systematically identify the dissipative channels associated with each dSIDM model.
[…]
Figure 1: Emission diagrams for dissipative DM self-scattering. The orange dots show other possible emission points. Depending on the type of diagram, we introduce the following terminology: (a) “External leg diagrams” for emission from external legs producing four diagrams for distinguishable and eight diagrams (including crossed channels) for identical particles. These diagrams are relevant for all considered models. (b) “Vertex diagrams” for scalar DM candidates $S$ and $\tilde{S}$, where emission occurs at the mediator-exchange vertex. They constitute two diagrams for distinguishable particles and, including crossed channels, four diagrams for identical particles. (c) “Contact diagrams” due to quartic scalar DM interactions; there are four diagrams of this type. (d) “Virtual internal Bremsstrahlung” (VIB) from a scalar mediator with scalar emission of $\phi$ particles, in principle relevant to all DM candidates; they lead to one (two) diagram(s) in the case of distinguishable (identical) particle scattering. The emitted particle $\phi$ or $V$ carries four-momentum $q$; the four-momentum assignment of propagators is emission-point dependent, and the full kinematic dependency is taken into account when evaluating the various amplitudes.
@@
We note that, in the particular case where a massless $V_{\mu}$ mediates the DM interaction and is also the emitted particle, the results are well-established in the literature, corresponding to QED- and scalar QED-like bremsstrahlung processes.
@@ 3.2 Non-relativistic hierarchy of scales
We consider scattering between DM candidates with a relative velocity that lies in the range $v_{i}\lesssim 10^{-2}$, corresponding to the low-velocity regimes discussed in Sec. 2.2. This NR assumption, when applied in the center-of-mass (CM) frame of the scattering processes, leads to a separation of scales between the relative incoming three-momentum, $\vec{p}_{i}\sim v_{i}\mu$, and the available CM-energy for the emission of a particle, $\omega\sim v_{i}^{2}\mu$$\mu$ is the reduced mass of the DM scattering pair.
@@
Futhermore, for the dSIDM scenarios involving scalar DM candidates:
@@
    Gauge vector emission: since the contact diagrams are velocity-suppressed, the leading order contributions for $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$ and $S_{1}S_{2}\to S_{1}S_{2}V$ come from the exchange diagrams (Fig. 1(a)), with the contributions from the vertex diagrams (Fig. 1(b)) being sub-leading. In the NR limit, spin effects from gauge interactions are suppressed [61]; see App. D for a detailed discussion. As a result, the amplitudes we derive are valid for general spin in the case of distinguishable particles. For identical DM particles, the Dirac fermions $\chi$ and the complex scalars $S$ must be treated separately since $t$- and $u$-channel amplitudes have differing relative sign. We define the following general coupling constant for these processes
@@
    Scalar emission from fermionic DM candidates: for the two dSIDM scenarios $\chi_{i}\chi_{j}\to\chi_{i}\chi_{j}\phi$ and $\tilde{\chi}_{i}\tilde{\chi}_{j}\to\tilde{\chi}_{i}\tilde{\chi}_{j}\phi$, the $v_{i}$-leading and sub-leading order contribution to their emission amplitudes come from the exchange diagrams (Fig.1(a)). This results in these candidates having the same emission amplitude in the case of distinguishable scattering. The same conclusion holds in the case of identical particle scattering because of the suppression of the annihilation channels. The relevance of VIB processes is treated separately in App. B. For the emission amplitudes corresponding to both models, we define
@@
    Scalar emission from scalar DM candidates: for the scenarios $S_{i}S_{j}\to S_{i}S_{j}\phi$ and $\tilde{S}_{i}\tilde{S}_{j}\to\tilde{S}_{i}\tilde{S}_{j}\phi$, the contact diagrams (Fig. 1(c)) and the ones mediated by $\tilde{S}$ are velocity suppressed. Consequently, the leading contributions to their emission amplitudes come from the exchange (Fig. 1(a)), and the vertex (Fig. 1(b)) diagrams. The relevance of VIB processes is treated separately in App. B. Therefore, except in specially constructed cases, both types of scalar DM candidates, $\tilde{S}$ and $S$, yield the same leading and sub-leading analytic expressions in a $v_{i}$-expansion of the overall emission amplitude. This result suggests a degree of model-independence also in the dissipative dynamics of scalar DM, allowing us to define
@@
In the opposite limit of a short-range mediated force, the emitted particles $(\phi,V^{\mu})$ necessarily differ from the exchanged particles $(\phi^{\prime},V^{\prime\mu},\tilde{S}_{j})$. This implies that, for the Lagrangian interactions discussed in Sec. 2, we must introduce a primed version of associated coupling constants $g^{\prime}_{j}$, $y^{\prime}_{j}$ and $A_{j}^{\prime}/m_{j}$. For $S_{1}S_{2}\to S_{1}S_{2}V$, this leads to the new interaction $Q_{j}Q_{j}^{\prime}g_{j}g_{j}^{\prime}V_{\mu}V^{\prime\mu}S_{j}^{\dagger}S_{j}$ corresponding to vertex emission (Fig.1(b)). [^5] This is a consequence of the covariant derivative being modified to $D_{\mu}=i\partial_{\mu}-igV_{\mu}-ig^{\prime}V^{\prime}_{\mu}$. Similarly, for scalar emission with scalar DM candidates introduced in (2.4), we account for the interactions $\lambda^{\prime}_{j}\phi\phi^{\prime}|S_{j}|^{2}$ and $\tilde{\lambda}^{\prime}_{j}/2\,\phi\phi^{\prime}\tilde{S}_{j}^{2}$. Finally, we consider a direct coupling between the light and heavy mediators only in the scalar-mediated, scalar-emission case, which gives rise to VIB processes (Fig. 1(d)). Since this contribution is significantly more suppressed than the other dissipative channels, we defer its discussion to App. B.2.
@@
Whereas the requirement for the presence of a contact interaction is mild, $m_{\phi^{\prime},V^{\prime}}\gtrsim v_{i}\mu\sim 10^{-3}\mu$, the typical coupling of $\phi^{\prime}$ or $V^{\prime}$ to DM must be larger than the coupling of $\phi$ or $V$ by a factor of $1/v_{i}\sim 10^{3}$. This demands a clear hierarchy of couplings in the dark sector. Such a scenario is then best conceivable in gauged models of SIDM where the DM self-interaction is due to (non-Abelian) strong interactions in the hidden sector.
The condition (3.2) also guides us to focus on mediator masses that satisfy, $m_{\phi^{\prime},V^{\prime}}<\sqrt{s}$ where $s=(p_{1}+p_{2})^{2}$ is the squared CM energy. Heavy mediators $m_{\phi^{\prime},V^{\prime}}>\sqrt{s}$ are, first, less prospective in successfully realizing SIDM scenarios with sufficiently large self-scattering cross-section—unless strong interactions are invoked—and second, typically do not dominate the dissipative process, thereby violating (3.2). Hence, by the standards of $\sqrt{s}\simeq m_{1}+m_{2}$ we continue to consider “light” mediators $\phi^{\prime}$ and $V^{\prime}$ as the prototype case, but such that they transmit a short-range force in the context of NR DM self-scattering. In this regime, we can then safely neglect any $s$-channel self-scattering diagrams that would otherwise contribute to the emission process. By the same token, we are also allowed to disregard contributions in scalar cases where the exchanged mediator is the DM itself.
@@
$\mathcal{M}_{\text{external leg}}+\mathcal{M}_{\text{contact}}$ $=\mathcal{F}(\lambda_{S},A_{j},A_{j}^{\prime},m_{j})\mathcal{T}(r,\omega,m_{\phi})+\mathcal{O}(r^{2}),$ (3.4)
@@
$\mathcal{M}_{\text{external leg}}+\mathcal{M}_{\text{vertex}}+\mathcal{M}_{\text{contact}}$ $=\mathcal{F}(\lambda_{S},g_{j},g_{j}^{\prime},m_{j})\mathcal{T}(r,\omega,m_{V})+\mathcal{O}(r^{2}).$ (3.5)
@@
In this section, we shall clarify the terminology of dipole and quadrupole emission used throughout this work, before presenting the formula for the energy differential cross section used in this work.
@@
$H_{0}=\frac{\vec{p}_{1}^{2}}{2m_{1}}+\frac{\vec{p}_{2}^{2}}{2m_{2}}+\frac{g_{1}g_{2}e^{-m_{V}|\vec{x}_{1}-\vec{x}_{2}|}}{|\vec{x}_{1}-\vec{x}_{2}|^{2}}.$ (3.6)
If the force is mediated by the scalar, the substitutions $g_{1}g_{2}\rightarrow y_{1}y_{2}$ and $m_{V}\rightarrow m_{\phi}$ apply in (3.6). The transition matrix element for the emission of a vector $V^{\mu}$ or scalar $\phi$ of energy $\omega$, three-momentum $\mathbf{q}$ (and of helicity $\lambda$ for $V^{\mu}$) is given by
@@
In the case of vector emission, we work in the Coulomb gauge, where the polarization vector $\epsilon=(0,\vec{e})$ reduces to its spatial component $\mathbf{e}$, which is transverse to the emitted vector momentum. Using the definition of the interaction potential given in (3.8a), the $\mathcal{T}$- matrix element up to the energy conserving delta function takes the form
@@
This result can be further simplified when the wavelength of the emitted particle is larger than the typical separation $|\mathbf{r}|$ of particles from the center of mass, i.e., $\nu\equiv\vec{r}\cdot\vec{q}\ll 1$. Expanding the exponential in (3.9) in $\nu$, and integrating over $d^{3}\mathbf{X}$ yields $\mathcal{T}_{fi}=(2\pi)^{4}\delta(E_{i}-E_{f}-\omega)\delta^{(3)}(\vec{p}_{i}+\vec{p}_{f}-\vec{q})\mathcal{M}$ where
@@
The first term describes electric dipole radiation and cancels in the case of equal charge-to-mass ratio. The second term describes electric quadrupole and magnetic dipole radiation. We note, however, that in a NR system, magnetic dipole effects can be neglected compared to electric ones [61]. This aligns with the classical expectation that the magnetic dipole moment of a two-particle system vanishes [96] [^6] Classically, the magnetic dipole momentum is proportional to the double derivative of the angular momentum of the system, which is conserved in a two particles system..
@@
In contrast to the vector emission, the $\mathcal{O}(\nu^{0})$ order vanishes, and (3.12) is composed of second- and third-order in $\nu$-terms. We, of course, encounter this structure again in the Feynman-diagrammatic calculations; see App. D. In analogy to the case of vector emission, the first term in (3.12), which cancels in the case of equal coupling-to-mass ratio, is identified as dipole radiation, and the second term as quadrupole radiation.
As can be seen from (3.10) and (3.12), quadrupole radiation is of higher order in velocity when compared to the dipole case: in the amplitudes by a factor of $\mathbf{r}\cdot\mathbf{q}\sim v_{i}$, and in terms of emission rates, by a factor of $v_{i}^{2}$. Therefore, dipole radiation dominates, unless it vanishes because of an identical charge-to-mass ratio of the colliding particle pair. In turn, this means that for identical particle scattering, the quadrupole order is the dominant emission mode.
@@
Energy-differential emission cross section
Having clarified the principal leading analytical structure of emission amplitudes, we close this section by quoting the NR formula for the energy differential $2\to 3$ emission cross section in the CM frame

### v2

@@ 3.1 Dissipative channels
We aim to derive the energy-differential cross section within a perturbative framework. For each dSIDM model, the emission amplitude $\mathcal{M}$ is constructed from Feynman diagrams as shown in Fig. 1. In the following, we systematically identify the dissipative channels associated with each dSIDM model.
11332244qqkk
11332244qqkk
33221144qq
11332244qqkqk-qkk
Figure 1: Emission diagrams for dissipative DM self-scattering. The orange dots show other possible emission points. Depending on the type of diagram, we introduce the following terminology: (a) “External leg diagrams” for emission from external legs producing four diagrams for distinguishable and eight diagrams (including crossed channels) for identical particles. These diagrams are relevant for all considered models. (b) “Vertex diagrams” for scalar DM candidates $S$ and $\tilde{S}$, where emission occurs at the mediator-exchange vertex. They constitute two diagrams for distinguishable particles and, including crossed channels, four diagrams for identical particles. (c) “Contact diagrams” due to quartic scalar DM interactions; there are four diagrams of this type. (d) “Virtual Internal Bremsstrahlung” (VIB) from a scalar mediator with scalar emission of $\phi$ particles, in principle relevant to all DM candidates; they lead to one (two) diagram(s) in the case of distinguishable (identical) particle scattering. The emitted particle $\phi$ or $V^{\mu}$ carries four-momentum $q$; the four-momentum assignment of propagators is emission-point dependent, and the full kinematic dependency is taken into account when evaluating the various amplitudes.
@@
We note that, in the particular case where a massless $V^{\mu}$ mediates the DM interaction and is also the emitted particle, the results are well-established in the literature, corresponding to QED- and scalar QED-like bremsstrahlung processes.
@@ 3.2 Non-relativistic hierarchy of scales
We consider scattering between DM candidates with a relative velocity that lies in the range $v_{i}\lesssim 10^{-2}$, corresponding to the low-velocity regimes discussed in Sec. 2.2. This NR assumption, when applied in the center-of-mass (CM) frame of the scattering processes, leads to a separation of scales between the relative incoming three-momentum, $\vec{p}_{i}\sim v_{i}\mu$, and the available CM-energy for the emission of a particle, $\omega\sim v_{i}^{2}\mu$; $\mu$ is the reduced mass of the DM scattering pair.
@@
Furthermore, for the dSIDM scenarios involving scalar DM candidates:
@@
    Gauge vector emission: since the contact diagrams are velocity-suppressed, the leading order contributions for $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$ and $S_{1}S_{2}\to S_{1}S_{2}V$ come from external leg diagrams (Fig. 1(a)), with the contributions from the vertex diagrams (Fig. 1(b)) being sub-leading. In the NR limit, spin effects from gauge interactions are suppressed [73]; see App. D for a detailed discussion. As a result, the amplitudes we derive are valid for general spin in the case of distinguishable particles. For identical DM particles, the Dirac fermions $\chi$ and the complex scalars $S$ must be treated separately since $t$- and $u$-channel amplitudes have differing relative sign. We define the following general coupling constant for these processes
@@
    Scalar emission from fermionic DM candidates: for the two dSIDM scenarios $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$ and $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, the $v_{i}$-leading and sub-leading order contribution to their emission amplitudes come from external leg diagrams (Fig.1(a)). This results in these candidates having the same emission amplitude in the case of distinguishable scattering. The same conclusion holds in the case of identical particle scattering because of the suppression of the annihilation channels. The relevance of VIB processes is treated separately in App. B. For the emission amplitudes corresponding to both models, we define
@@
    Scalar emission from scalar DM candidates: for the scenarios $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$, the contact diagrams (Fig. 1(c)) and the ones mediated by $\tilde{S}$ are velocity suppressed. Consequently, the leading contributions to their emission amplitudes come from external leg (Fig. 1(a)), and vertex (Fig. 1(b)) diagrams. The relevance of VIB processes is treated separately in App. B. Therefore, except in specially constructed cases, both types of scalar DM candidates, $\tilde{S}$ and $S$, yield the same leading and sub-leading analytic expressions in a $v_{i}$-expansion of the overall emission amplitude. This result suggests a degree of model-independence also in the dissipative dynamics of scalar DM, allowing us to define
@@
In the opposite limit of a short-range mediated force, the emitted particles $(\phi,V^{\mu})$ necessarily differ from the exchanged particles $(\phi^{\prime},V^{\prime\mu},\tilde{S}_{j})$. This implies that, for the Lagrangian interactions discussed in Sec. 2, we must introduce a primed version of associated coupling constants $g^{\prime}_{j}$, $y^{\prime}_{j}$ and $A_{j}^{\prime}/m_{j}$. For $S_{1}S_{2}\to S_{1}S_{2}V$, this leads to the new interaction $g_{j}g_{j}^{\prime}V_{\mu}V^{\prime\mu}S_{j}^{\dagger}S_{j}$ corresponding to vertex emission (Fig.1(b)). [^5] This is a consequence of the covariant derivative being modified to $D_{\mu}=i\partial_{\mu}+igV_{\mu}+ig^{\prime}V^{\prime}_{\mu}$. Similarly, for scalar emission with scalar DM candidates introduced in (2.4), we account for the interactions $\lambda^{\prime}_{j}\phi\phi^{\prime}|S_{j}|^{2}$ and $\tilde{\lambda}^{\prime}_{j}/2\,\phi\phi^{\prime}\tilde{S}_{j}^{2}$. Finally, we consider a direct coupling between the light and heavy mediators only in the scalar-mediated, scalar-emission case, which gives rise to VIB processes (Fig. 1(d)). Since this contribution is significantly more suppressed than the other dissipative channels, we defer its discussion to App. B.2.
@@
Whereas the requirement for the presence of a contact interaction is mild, $m_{\phi^{\prime},V^{\prime}}\gtrsim v_{i}\mu\sim 10^{-3}\mu$, the typical coupling of $\phi^{\prime}$ or $V^{\prime\mu}$ to DM must be larger than the coupling of $\phi$ or $V^{\mu}$ by a factor of $1/v_{i}\sim 10^{3}$. This demands a clear hierarchy of couplings in the dark sector. Such a scenario is then best conceivable in gauged models of SIDM where the DM self-interaction is due to (non-Abelian) strong interactions in the hidden sector.
The conditions (3.2) and (3.3) also guide us to focus on mediator masses that satisfy, $m_{\phi^{\prime},V^{\prime}}<\sqrt{s}$ where $s=(p_{1}+p_{2})^{2}$ is the squared CM energy. Heavy mediators $m_{\phi^{\prime},V^{\prime}}>\sqrt{s}$ are, first, less prospective in successfully realizing SIDM scenarios with sufficiently large self-scattering cross section—unless strong interactions are invoked—and second, typically do not dominate the dissipative process, thereby violating (3.2)–(3.3). Hence, by the standards of $\sqrt{s}\simeq m_{1}+m_{2}$ we continue to consider “light” mediators $\phi^{\prime}$ and $V^{\prime\mu}$ as the prototype case, but such that they transmit a short-range force in the context of NR DM self-scattering. In this regime, we can then safely neglect any $s$-channel self-scattering diagrams that would otherwise contribute to the emission process. By the same token, we are also allowed to disregard contributions in scalar cases where the exchanged mediator is the DM itself.
@@
$\mathcal{M}_{\text{external leg}}+\mathcal{M}_{\text{contact}}$ $=\mathcal{F}(\lambda_{S},A_{j},A_{j}^{\prime},m_{j})\mathcal{T}(r,\omega,m_{\phi})+\mathcal{O}(r),$ (3.4)
@@
$\mathcal{M}_{\text{external leg}}+\mathcal{M}_{\text{vertex}}+\mathcal{M}_{\text{contact}}$ $=\mathcal{F}(\lambda_{S},g_{j},g_{j}^{\prime},m_{j})\mathcal{T}(r,\omega,m_{V})+\mathcal{O}(r).$ (3.5)
@@
In this section, we shall clarify the terminology of dipole and quadrupole emission used throughout this work, before presenting the formula for the energy-differential cross section.
@@
$H_{0}=\frac{\vec{p}_{1}^{2}}{2m_{1}}+\frac{\vec{p}_{2}^{2}}{2m_{2}}+\frac{g_{1}g_{2}e^{-m_{V}|\vec{x}_{1}-\vec{x}_{2}|}}{4\pi|\vec{x}_{1}-\vec{x}_{2}|^{2}}.$ (3.6)
If the force is mediated by the scalar, the substitutions $g_{1}g_{2}\rightarrow-y_{1}y_{2}$ and $m_{V}\rightarrow m_{\phi}$ apply in (3.6). The transition matrix element for the emission of a vector $V^{\mu}$ or scalar $\phi$ of energy $\omega$, three-momentum $\mathbf{q}$ (and of helicity $\lambda$ for $V^{\mu}$) is given by
@@
In the case of vector emission, we work in the Coulomb gauge, where the polarization vector $\varepsilon^{\mu*}(q)=(0,\vec{e}^{*})$ reduces to its spatial component $\mathbf{e}^{*}$, which is transverse to the emitted vector momentum. Using the definition of the interaction potential given in (3.8a), the $\mathcal{T}$- matrix element, up to the energy conserving delta function and the factor $2\pi i$, takes the form
@@
This result can be further simplified when the wavelength of the emitted particle is larger than the typical separation $|\mathbf{r}|$ of particles from the center of mass, i.e., $\nu\equiv\vec{r}\cdot\vec{q}\ll 1$. Expanding the exponential in (3.9) in $\nu$, and integrating over $d^{3}\mathbf{X}$ yields $\mathcal{T}_{fi}=-i(2\pi)^{4}\delta(E_{i}-E_{f}-\omega)\delta^{(3)}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4}-\vec{q})\mathcal{M}$ where
@@
In our setting, i.e., for a two-particles system without explicit magnetic moment interaction, for massless vector particles the second term in (3.10) corresponds to electric quadrupole radiation. [^7] If spin effects are included, the magnetic moment yields an additional interaction potential $V_{\text{int}}^{\prime}=-\sum_{n}\bm{\mu}_{n}\cdot\left(\text{curl}\,\mathbf{A}(\vec{x}_{n})\right)$, where $|\bm{\mu}_{n}|=g_{n}/(2m_{n})$ for a $1/2$-spin particle. In a NR system, however, the latter can be neglected compared to electric ones both at dipole and quadrupole order [73]. Using transversality of the polarization state and the canonical commutation relations, the operator in the bracket can be reduced to its symmetric part,
[…]
Squaring the corresponding matrix element, summing over the transverse polarization states and averaging over the emitted particle direction $\bm{\hat{}}{\mathbf{q}}$ yields
[…]
where the subscript $\mathcal{Q}$ labels the quadrupole nature of the emission. The angular average decomposes into symmetric and anti-symmetric parts,
$\Big\langle\sum_{\lambda}{\hat{q}}^{j}{\hat{q}}^{k}{e}_{(\lambda)}^{*l}{e}_{(\lambda)}^{m}\Big\rangle_{\bm{\hat{}}{\mathbf{q}}}$ $=\frac{1}{15}\left[\frac{5}{2}\left(\delta^{jk}\,\delta^{lm}-\delta^{jm}\,\delta^{kl}\right)+\frac{3}{2}\left(-\frac{2}{3}\,\delta^{jl}\,\delta^{km}+\delta^{jm}\,\delta^{kl}+\delta^{jk}\,\delta^{lm}\right)\right],$ (3.13)
respectively. The first, antisymmetric part cancels in the contraction (3.12), yielding the familiar, trace-free Cartesian quadrupole form,
[…]
We note in passing, that the absence of magnetic dipole contributions is also in agreement with the classical expectation of vanishing magnetic dipole radiation for a two-particle system [103].
@@
In contrast to the vector emission, the $\mathcal{O}(\nu^{0})$ order vanishes, and (3.16) is composed of second- and third-order-in-$\nu$ terms. We, of course, encounter this structure again in the Feynman-diagrammatic calculations; see App. D. In analogy to the case of vector emission, the first term in (3.16), which cancels in the case of equal coupling-to-mass ratio, is identified as a dipole transition.
By squaring and averaging the second term over the direction of the emitted particle, one obtains a contraction with
$\left\langle{\hat{q}}^{j}{\hat{q}}^{k}{\hat{q}}^{l}{\hat{q}}^{m}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}=\frac{1}{15}\left(\delta^{jm}\delta^{kl}+\delta^{jk}\delta^{lm}+\delta^{jl}\delta^{km}\right).$ (3.17)
The second term of (3.16) then yields the contribution
@@
which, in contrast to (3.14), includes a trace component. Throughout the paper, we still refer to such scalar emission as being “quadrupole”, having in mind the relative velocity ordering rather than a specific tensorial structure. Indeed, as can be seen from (3.10) and (3.16), quadrupole radiation is of higher order in velocity when compared to the dipole case: in the amplitudes by a factor of $\mathbf{r}\cdot\mathbf{q}\sim v_{i}$, and in terms of emission rates, by a factor of $v_{i}^{2}$. Therefore, dipole radiation dominates, unless it vanishes because of an identical charge-to-mass ratio of the colliding particle pair. In turn, this means that for identical particle scattering, the quadrupole order is the dominant emission mode. At this point, we note that the Hamiltonian in (3.8b)—on which (3.18) is based on—receives model-dependent corrections at quadrupole order; the expressions derived below take this into account.
[…]
Energy-differential cross section
Having clarified the principal leading analytical structure of emission amplitudes, we close this section by quoting the NR formula for the energy-differential cross section in the CM frame